OpenStudy (anonymous):

The table shows the results of drawing letter tiles from a bag. What is the probability that the next tile drawn will have the letter A on it? A. 3/25 B.6/25 C.12/25 D. 3/5

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

@texaschic101

OpenStudy (mathmale):

You'll need to post the table you've mentioned.

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (texaschic101):

I am not good with probabilities....but I think you add up the tiles...(12 + 5 + 18 + 15) = 50. Then you put the A tiles over the total tiles, which is 12/50....then reduce it , and it becomes 3/5. You better wait for someone who knows exactly what they are doing, because I am not 100% sure this is correct.

OpenStudy (mathmale):

First of all, how many tiles, total, are in that bag? Secondly, how many of them are A tiles? sorry to be repeating the info given to you by texaschic101, but these are important questions that you need to learn to ask in this type of problem.

OpenStudy (texaschic101):

go ahead and explain mathmale...because I am not the best at probabilities

OpenStudy (anonymous):

IT DOESNT SAY HOW MAY TILES THEIR ARE @mathmale

OpenStudy (mathmale):

that means we bring out our adding machines (calculators) and add up all of the counts of the different lettr tiles.

OpenStudy (anonymous):

OK WHAT ARE THE COUNTS OF THE DIFFERENT LETTER TILES?

OpenStudy (mathmale):

12+5+18+15=??

OpenStudy (mathmale):

Please refer to the table.

OpenStudy (anonymous):

That equals 50 added together

OpenStudy (mathmale):

Yes. Notice that texaschic has that "50" in her solution? OK. How many of those tiles are A tiles?

OpenStudy (anonymous):

12 are A tiles?

OpenStudy (mathmale):

Yes. So the fraction of the total number of tiles in that bag that happen to be A tiles is 12/50. This is ALSO the probability that you were looking for. If you start with all 50 tiles in the bag and draw one at random, the chances are 12 in 50 that you will draw an A tile.

OpenStudy (mathmale):

Now please reduce that fraction. 12/50 = what fraction, in lowest terms?

OpenStudy (anonymous):

is it 6/25

OpenStudy (mathmale):

Yes, it is. Can't reduce that any further, so yes, the prob. tht you will draw an A tile is 6/25, or 0.24. Either form of the answer is fine.

OpenStudy (texaschic101):

I am so sorry...I messed up on the simple reducing. Thanks mathmale for doing this....she would have gotten the wrong answer if she listened to me.